OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation y - 4x = -3

jimthompson5910 (jim_thompson5910):

solve for y to get y - 4x = -3 y - 4x+4x = -3+4x y + 0x = -3+4x y = 4x-3 now compare it to y = mx+b

jimthompson5910 (jim_thompson5910):

remember that y = mx+b is the general slope intercept form where m = slope b = y-intercept

OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. @jim_thompson5910

OpenStudy (anonymous):

put the equation into \[y=mx+b\] then replace y with 0 to get your first x-intercept and then replace x with 0 to get your y-intercept

jimthompson5910 (jim_thompson5910):

to find the y-intercept, plug in x = 0 and solve for y

jimthompson5910 (jim_thompson5910):

to find the x-intercept, plug in y = 0 and solve for x